OpenStudy (anonymous):

Solve: −2(x − 3) = −6 0 −6 6 −12

OpenStudy (nurali):

−2(x − 3) = −6 -2x+6=-6 -2x=-6-6 -2x=-12 2x=12 x=12/2 x=6

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

yup thats correct x=6

OpenStudy (anonymous):

np